# Marine spatial planning techniques with a case study on wave-powered offshore aquaculture farms

### v2

@@ Abstract
As emerging marine technologies lead to the development of new infrastructure across the ocean, they enter an environment that existing ecosystems and industries already rely on. Although necessary to provide sustainable sources of energy and food, careful planning will be important to make informed decisions and avoid conflicts. This paper examines several techniques used for marine spatial planning, an approach for analyzing and planning the use of marine resources. Using open source software including QGIS and Python, the potential for developing wave-powered offshore aquaculture farms using the RM3 wave energy converter along the Northeast coast of the United States is assessed and several feasible sites are identified. The optimal site, located at 43.7°N 68.9°W along the coast of Maine, has a total cost for a 5-pen farm of $56.8M, annual fish yield of 676 tonnes, and a levelized cost of fish of $9.23 per kilogram. Overall trends indicate that the cost greatly decreases with distance to shore due to the greater availability of wave energy and that conflicts and environmental constraints significantly limit the number of feasible sites in this region.
@@ 1 Introduction
Oceans have long been a source of opportunity, and conflict, for the communities and industries along their shores. This rich environment, first used as a source of food and for recreation, has since been utilized by other offshore activities including mineral extraction, shipping, aquaculture, and most recently, offshore renewable energy (ORE) systems. Even before the expansion of renewable energy, conflicts between different industries led to challenging situations between fisheries and other stakeholders with varying degrees of resolution. Arbo and Thủy (2016), for example, researched conflicts between the oil industry and fisheries in Norway and Vietnam. In Norway, the two industries were able to reach a common agreement to collaborate and work together with an integrated management plan in the Barents Sea, while in Vietnam the two industries have continued to face challenging conflicts. Another study by Andrews et al. (2021) examined the offshore oil industry’s adverse impacts on the environment, small-scale fisheries, and coastal community livelihood. Conflicts involving fisheries are further complicated by uncontrolled fishing activities and additional conflicts between recreational and commercial fisheries, which threaten fish stocks and the overall health of the ocean (Bess and Rallapudi, 2007).
@@
Entering this mix is the emerging industry of offshore aquaculture, which has the potential to provide a more sustainable and protein-rich source of food for the growing population. Compared to conventional aquaculture, offshore aquaculture farms raise fish in the open ocean instead of sheltered bays or onshore tanks. Existing literature by researchers including Gentry et al. (2017), Weiss et al. (2018), Brugere (2006), and Bishwajit (2014) has brought to light challenges and opportunities for aquaculture farms, including the potential for significant expansion in many parts of the world including Southeast Asia, Australia, North America, and South America. Many of these studies, and particularly Longdill et al. (2008), look specifically at the impact on the local environment and how different environmental factors and the siting of farms can change the effects of pollution and other negative disturbances. However, few consider the energy source of the farm, which is often diesel generators, and the potential ways to mitigate the environmental impacts of the associated fossil fuel use. This paper aims to assess the potential to integrate wave power, a renewable source of energy, into the farm setup, in addition to analyzing the potential conflicts and environmental constraints that such a farm might face.
@@
High ORE growth in response to climate change has created new challenges for many fisheries that currently use the same waters. Stelzenmüller et al. (2022), studied the impacts of ORE development in the North Sea and the socioeconomic effects on fisheries in that region. Although they find that conflicts exist currently and are ongoing, they predict increasing long-term conflicts between fisheries and ORE activities after 2025. Some work does show promising signs for the collaboration between existing fisheries and the expanding ORE industry, particularly when coastal communities are brought into the planning process early on. The Block Island Wind Farm, for example, was able to work with local communities and fisheries and decided to change certain plans, including the placement of the turbines, to minimize conflict in response to community input (Firestone et al., 2020; ten Brink and Dalton, 2018).
@@
Using these techniques, this paper aims to assess and provide suggestions for the development of a sustainable offshore aquaculture farm powered by wave energy in the Northeast United States. By beginning with large-scale mapping techniques, promising sites in the region of interest are identified, and details of the overall considerations that must be made when co-locating wave energy converters and aquaculture farms are revealed. Further integration with Python allows GIS information to be used in a more complex modeling process that includes an objective function to identify cost-effective sites while meeting several constraints. These techniques exemplify the ways that GIS information and modeling could be integrated into a wide range of MSP processes to improve the understanding of the oceanic environment, pick out overall trends relevant to a project, and initiate early engagement with stakeholders to avoid or mitigate potential conflicts.
@@ 2 Problem formulation
The modeled WPAF includes five cylindrical net pens of 30 m diameter and 15 m height, each with a fish stocking density of . The farm will be stocked with Atlantic salmon, a common and in-demand species. The WECs used to power the farm will be the Reference Model 3 (RM3) as designed by Sandia National Laboratories. The RM3 WEC is a two-body heaving point absorber with a top float that moves relative to a vertical column and reaction place when excited by waves. This mechanical energy from ocean waves is converted into electrical energy through a power take-off (PTO) system and a generator within the WEC. The RM3 has a height of 30 m, float diameter of 20m, and a rated capacity of 286 kW (Neary et al., 2014). Per discussion with aquaculture stakeholders, a spacing of 150 m between the net pens is considered to provide enough space for the mooring and vessel operation. Additionally, the spacing between WECs is considered twice the WEC diameter to avoid any destructive wave interaction and reduction in the generated power (Zhong and Yeung, 2019). The images in figure 2 show a submersible net pen and RM3 WEC like what is considered in this study. Two feed barges are also required to enable automatic fish feeding; each barge has six feed lines and a capacity of 200 metric tons (Scale Aquaculture AS, 2019). Travel to the net pens using ships is scheduled weekly to restock the feed barges and check the aquaculture farm.
@@ 3.1 Project scope
There are several approaches to evaluating complex models across a large region, depending on the needs of a project. Projects aiming to identify individual optimal sites may benefit from optimization techniques such as gradient descent that search for locations that minimize an objective function such as cost. For this project, however, a brute-force approach is used to allow the entire region to be evaluated so that overall trends can easily be identified. Unlike an optimization approach which strategically searches for a minimum, the brute-force approach evaluates every location at a set interval across the region. Although more time-consuming, the results of this approach provides a more complete understanding of the entire region. The results using this approach are achieved in a reasonable time frame, under and hour, for the current model, however a more complex model may necessitate a different approach to optimization. The interval at which sites are evaluated with this approach determines the “resolution” of the results, which is limited by how much processing time is practical and the lowest resolution of the data sets used in the project. Some common properties like bathymetry have very high-resolution data sets available, while others may be lower resolution or limited to smaller areas. For this project, a resolution of 0.1° longitude and latitude (roughly 10km in this region) is selected to balance the need for precision while not being overly demanding of computational time. This is slightly lower than the most limiting data sets of wave conditions, which have a resolution of about 0.067°. Figure 3 shows the points sampled at this resolution alongside the state and federal waters included in the project scope. While this resolution provides general trends across the entire region, a higher resolution pass would be necessary for later stages of planning, especially once a more specific region of interest can be identified. Table 1 includes the resolution of individual data files, which indicate the highest resolution that could be achieved before other data sources become necessary.
@@ 3.2 Sources of data
GIS data types fall into two general types: raster and vector data sets. Raster data sets provide pixels of continuous information like physical conditions or satellite imagery, while vector data sets describe points, lines, or areas. In this study, raster data sets are used for oceanic conditions and bathymetry, while vector data sets describe regions of potential conflict and administrative boundaries. Current speed, dissolved oxygen, temperature, and salinity are used to determine the carrying capacity. Significant wave height and wave energy period are used to determine the wave power potential, which along with distance to port is used in the cost model. Bathymetry data and regions of conflicting industries and restricted areas are used to constrain the feasible area. Together, these data sets are used to evaluate the location-dependent costs for each site and to identify potential conflicts with other industries and restricted areas, as discussed in sections 5.2 and 5.3.
@@
It is important to select data that is a good fit for the project, especially when multiple sources are available. With significant wave height and wave period, for example, an older raster data set of annual averages from NREL is chosen because it provides continuous information across the region and is available in a format that is readily compatible with GIS. NOAA’s National Data Buoy Center, another possible option, could have provided the same information but would have limited the analysis to the few locations where buoys already exist and would have required more processing to integrate into the model. In other cases, available data sets may be limited by resolution or region. An older bathymetry data set, for example, is selected because it provides a high resolution across the entire project scope while more recent surveys are limited to a smaller region.
@@ 3.3 Software selection
Software is selected for this work with the intention of being accessible to a wide range of applications and users who might benefit from MSP. As a result, only open source software is used in the final process. GIS software is used for basic processing and visualization of results including the layout of published maps. QGIS is used as the main GIS software because of its widespread availability and complete set of processing tools. QGIS unfortunately lacks the ability to process certain proprietary data types from ArcGIS, the current industry standard, but is otherwise a capable and free option. Python, along with the GeoPandas package, is used to handle more complex modeling such as the carrying capacity and cost functions which would be difficult in GIS software alone. For these cases, a Python package is created to interface between the main model and the individual GIS data sets, as demonstrated in section 5.
@@ 4.1 Preprocessing data
The following preprocessing steps are necessary to standardize downloaded GIS data files and prepare them for further mapping and analysis. General considerations for each step are presented in this section, while instructions to complete them in QGIS are provided in A.
4.1.1 File conversion
The first step, conversion from an ArcGIS format, is only necessary for certain raster files distributed in the proprietary GeoDatabase (.gdb) format. Although QGIS can currently access vector data in these files, rasters must be loaded into ArcGIS and exported as GeoTIFF (.tiff) files to be readable by QGIS and other applications. This process is straightforward but unfortunately requires users interested in using data from organizations that publish data in GeoDatabases to have access to ArcGIS.
4.1.2 Vector interpolation
Certain data sets for environmental conditions are distributed in a vector format with many points containing measurements instead of as a continuous raster. For many analyses, having a continuous raster of conditions is beneficial, and interpolation can be used to approximate this from discrete measurements. For this study, the data used for dissolved oxygen and salinity are distributed as vector files, but are converted to rasters using this method.
4.1.3 Clipping to the project scope
The data is clipped to only include the scope of designated state and federal waters. This relies on a vector layer describing the project scope which, in this case, is the file with state and federal waters from NOAA in table 2.
4.1.4 Re-projection
GIS data and maps use different projections to translate the three-dimensional globe into a two-dimensional image. Many projections are available to balance the diversity of needs that a map might have, such as correctly displaying geometry at a certain latitude or ensuring that the area of shapes in different parts of a map appear to be the same size. Two-dimensional maps of large areas generally cannot accurately show every characteristic, so projections cater to the few that are most important for a given application. This project uses WGS 84 (EPSG:4326) and WGS 84 / UTM Zone 19 (EPSG:32619) for most steps. WGS 84 is a very common CRS that maintains parallel lines of longitude but causes distortion at higher latitudes. WGS 84 is used to standardize all downloaded data because of its simplicity, and so that each data set uses units of degrees in latitude and longitude. The Python tool, described in section 5, requires different files to have consistent units, and degrees are convenient for querying specific points in the ocean. WGS 84 / UTM Zone 19 is used for final maps exported from QGIS. This projection is tailored for the specific region of the earth included in the scope and minimizes distortion in that area. As a good fit for the Northeast United States, UTM Zone 19 allows published maps to look more visually accurate. Figure 4 shows a comparison of these two projections; notice how the angle of lines of latitude and longitude and the shapes of states vary between the two.
@@ 4.2 Raster analysis
Although most analysis for this case study is done within the Python model shown in section 5, there are several helpful functions in QGIS that can aid in understanding and visualizing the data. For this study, two are used: generating contours to display the range of oceanic conditions and finding regions that meet certain conditions such as the biological needs of the target species or the absence of conflicts with other activities. Figure 5 shows all raster data sets used in the case study, including environmental conditions used to identify favorable conditions for a certain species, average wave period and height to estimate WEC power output, and bathymetry and distance to port which further constrain feasible sites and affect the cost. The environmental conditions are also used to evaluate the carrying capacity of a location, which indicates how large a population can be sustained in a given environment. Contours, as discussed below, are overlaid on the raster layers.
Contours are simple to generate in QGIS and help to visualize and give a numerical reference for how conditions vary across the region. In the case of temperature, for example, figure 5 shows a clear north-south variation indicating that a species like Atlantic salmon, which prefer colder conditions, may be better suited for areas north of Cape Cod. The legend and raster alone give a sense of the overall range, such as the minimum average temperature of 7.5°C in this region. But the contours give more specificity such as the higher variation in southern regions and exact isolines of certain values of interest.
The Raster Calculator is another powerful QGIS tool with many uses, including calculating simple algebraic functions and filtering regions that meet certain conditions. Although these uses are handled in Python for this project, the Raster Calculator could be a helpful tool for similar analyses being completed entirely within QGIS. In this project, Raster Calculator is used to reclassify data on fishing traffic and convert it to a vector that can be used in the conflict analysis done to identify overlapping activities. Fishing data, which is distributed as a raster showing the annual number of vessels passing through a given pixel of the ocean, is first reclassified by the standard deviation of 20.3 around the mean of 4.6 vessels before being converted into a vector data set that forms a constraint.
@@ 4.3 Vector analysis
Vector analysis involves data sets made up of points, lines, or areas, rather than continuous values. In this study vector data is used to represent conflicts, which QGIS can aid in identifying. It is often helpful to convert between raster and vector formats depending on the analysis needed. Fishing vessel traffic, for example, is considered a conflict to be excluded in this analysis and so a vector of regions that have high fishing traffic is first created using the raster input data. Vector data sets are the core of the conflict analysis used to identify overlapping industries, and each of the included data sets are pictured in figure 6.
Like raster files, these data sets are loaded into Python later for this case study. However, preliminary maps of conflicts can also be generated in QGIS to identify and highlight areas of potential conflict early on. Figure 7 demonstrates this technique using the data sets for wind lease areas and regions of high fishing traffic, and highlights the regions where these industries overlap and therefore could conflict. Note that although the regions highlighted in figure 7 include conflicts with areas that had above-average fishing traffic in 2021, only areas with very high fishing traffic (over one standard deviation above the average) are considered infeasible farm locations in the Python model. This conclusion was determined after attending meetings hosted by the Bureau of Ocean Energy Management with stakeholders in the fishing and offshore wind industries. The fishing community is very protective of their fishing grounds so although available information is limited, but this assumption reflects a reasonable understanding of potential conflicts with fisheries and other industries.

### v3

@@ Abstract
As emerging marine technologies lead to the development of new infrastructure across the ocean, they enter an environment that existing ecosystems and industries already rely on. Although necessary to provide sustainable sources of energy and food, careful planning will be important to make informed decisions and avoid conflicts. This paper examines several techniques used for marine spatial planning, an approach for analyzing and planning the use of marine resources. Using open-source software including QGIS and Python, the potential for developing offshore aquaculture farms powered by a reference model wave energy converter from the Sandia National Labs, the RM3, along the Northeast coast of the United States is assessed and several feasible sites are identified. The optimal site, located at 43.7°N 68.9°W along the coast of Maine, has a total cost for a 5-pen farm of $56.8M, annual fish yield of 676 tonnes, and a levelized cost of fish of $9.23 per kilogram. Overall trends indicate that the cost greatly decreases with distance to shore due to the greater availability of wave energy and that conflicts and environmental constraints significantly limit the number of feasible sites in this region.
@@ 1 Introduction
Oceans have long been a source of opportunity, and conflict, for the communities and industries along their shores. This rich environment, first used as a source of food and for recreation, has since been utilized by other offshore activities including mineral extraction, shipping, aquaculture, and most recently, offshore renewable energy (ORE) systems. Even before the expansion of renewable energy, interactions between different industries led to challenging situations between fisheries and other stakeholders with varying degrees of resolution. Arbo and Thủy (2016), for example, researched disputes between the oil industry and fisheries in Norway and Vietnam. In Norway, the two industries were able to reach a common agreement to collaborate and work together with an integrated management plan in the Barents Sea, while in Vietnam the two industries have continued to face challenging conflicts. Another study by Andrews et al. (2021) examined the offshore oil industry’s adverse impacts on the environment, small-scale fisheries, and coastal community livelihood. Conflicts involving fisheries are further complicated by uncontrolled fishing activities and overlap between recreational and commercial fisheries, which threaten fish stocks and the overall health of the ocean (Bess and Rallapudi, 2007).
@@
Entering this mix is the emerging industry of offshore aquaculture, which has the potential to provide a more sustainable and protein-rich source of food for the growing population (NOAA Fisheries, 2023). Compared to conventional aquaculture, offshore aquaculture farms raise fish in the open ocean instead of sheltered bays or onshore tanks. Existing literature by researchers including Gentry et al. (2017), Weiss et al. (2018), Brugere (2006), and Bishwajit (2014) has brought to light challenges and opportunities for aquaculture farms, including the potential for significant expansion in many parts of the world including Southeast Asia, Australia, North America, and South America. Many of these studies, and particularly Longdill et al. (2008), look specifically at the impact on the local environment and how different environmental factors and the siting of farms can change the effects of pollution and other negative disturbances. However, few consider the energy source of the farm, which is often diesel generators, and the potential ways to mitigate the environmental impacts of the associated fossil fuel use.
@@
High ORE growth in response to climate change has created new challenges for many fisheries that currently use the same waters. Stelzenmüller et al. (2022), studied the impacts of ORE development in the North Sea and the socioeconomic effects on fisheries in that region. Although they find that conflicts exist currently and are ongoing, they predict increasing friction between fisheries and ORE activities after 2025. Some work does show promising signs for the collaboration between existing fisheries and the expanding ORE industry, particularly when coastal communities are brought into the planning process early on. The Block Island Wind Farm, for example, was able to work with local communities and fisheries and decided to change certain plans, including the placement of the turbines, in response to community input (Firestone et al., 2020; ten Brink and Dalton, 2018).
@@
Using these techniques, the authors of this paper aim to assess the potential to integrate wave power into the farm setup, analyze the potential conflicts and environmental constraints that such a farm might face, and provide suggestions for the development of such a farm in the Northeast United States. The farm is modeled using Atlantic salmon, the reference model three (RM3) WEC, and submersible net pens like those from Innovasea, as explained in section 2. By beginning with large-scale mapping techniques, promising sites in the region of interest are identified, and details of the overall considerations that must be made when co-locating wave energy converters and aquaculture farms are revealed. Further integration with Python allows GIS information to be used in a more complex modeling process that includes an objective function to identify cost-effective sites while meeting several constraints. These techniques exemplify the ways that GIS information and modeling could be integrated into a wide range of MSP processes to improve the understanding of the oceanic environment, pick out overall trends relevant to a project, and initiate early engagement with stakeholders to avoid or mitigate potential issues.
@@ 2 Problem formulation
The modeled WPAF includes five cylindrical net pens of 30 m diameter and 15 m height, each with a fish stocking density of . The net pen design is similar to those designed by Innovasea, and will be stocked with Atlantic salmon, a common and in-demand species with similar biological needs to steelhead trout, as the target species. The WECs used to power the farm will be the Reference Model 3 (RM3) as designed by Sandia National Laboratories. The RM3 WEC is a two-body heaving point absorber with a top float that moves relative to a vertical column and reaction place when excited by waves. This mechanical energy from ocean waves is converted into electrical energy through a power take-off (PTO) system and a generator within the WEC. The RM3 has a height of 30 m, float diameter of 20m, and a rated capacity of 286 kW (Neary et al., 2014). Per discussion with aquaculture stakeholders, a spacing of 150 m between the net pens is considered to provide enough space for the mooring and vessel operation. Additionally, the spacing between WECs is considered twice the WEC diameter to avoid any destructive wave interaction and reduction in the generated power (Zhong and Yeung, 2019). Two feed barges are also required to enable automatic fish feeding; each barge has six feed lines and a capacity of 200 metric tons (Scale Aquaculture AS, 2019). Travel to the net pens using ships is scheduled weekly to restock the feed barges and check the aquaculture farm.
@@ 3.1 Study scope
There are several approaches to evaluating complex models across a large region, depending on the needs of a project. Projects aiming to identify individual optimal sites may benefit from optimization techniques such as gradient descent that search for locations that minimize an objective function such as cost. For this study, however, a brute-force approach is used to allow the entire region to be evaluated so that overall trends can easily be identified. Unlike an optimization approach which strategically searches for a minimum, the brute-force approach evaluates every location at a set interval across the region. Although more time-consuming, the results of this approach provides a more complete understanding of the entire region. The results using this approach are achieved in a reasonable time frame, under an hour, for the current model, however a more complex model may necessitate a different approach to optimization. The interval at which sites are evaluated with this approach determines the “resolution” of the results, which is limited by how much processing time is practical and the lowest resolution of the data sets used in the study. Some common properties like bathymetry have very high-resolution data sets available, while others may be lower resolution or limited to smaller areas. For this study, a resolution of 0.1° longitude and latitude (roughly 10km in this region) is selected to balance the need for precision while not being overly demanding of computational time. This is slightly lower than the most limiting data sets of wave conditions, which have a resolution of about 0.067°. Figure 3 shows the points sampled at this resolution alongside the state and federal waters included in the study scope. While this resolution provides general trends across the entire region, a higher resolution pass would be necessary for later stages of planning, especially once a more specific region of interest can be identified. Table 1 includes the resolution of individual data files, which indicate the highest resolution that could be achieved before other data sources become necessary.
@@ 3.2 Sources of data
GIS data types fall into two general types: raster and vector data sets. Raster data sets provide pixels of continuous information like physical conditions or satellite imagery, while vector data sets describe points, lines, or areas. In this study, raster data sets are used for oceanic conditions and bathymetry, while vector data sets describe regions of potential conflict and administrative boundaries. Current speed, dissolved oxygen, temperature, and salinity are used to determine the carrying capacity. Significant wave height and wave energy period are used to determine the wave power potential, which along with distance to port is used in the cost model. Bathymetry data, regions with other active industries, and restricted areas are used to constrain the feasible area. Together, these data sets are used to evaluate the location-dependent costs for each site and identify potential conflicts, as discussed in sections 5.2 and 5.3.
@@
It is important to select data that is a good fit for the study, especially when multiple sources are available. With significant wave height and wave period, for example, an older raster data set of annual averages from NREL is chosen because it provides continuous information across the region and is available in a format that is readily compatible with GIS. NOAA’s National Data Buoy Center, another possible option, could have provided the same information but would have limited the analysis to the few locations where buoys already exist and would have required more processing to integrate into the model. In other cases, available data sets may be limited by resolution or region. An older bathymetry data set, for example, is selected because it provides a high resolution across the entire study scope while more recent surveys are limited to a smaller region.
@@ 3.3 Software selection
Software is selected for this work with the intention of being accessible to a wide range of applications and users who might benefit from MSP. As a result, only open-source software is used in the final process. GIS software is used for basic processing and visualization of results including the layout of published maps. QGIS is used as the main GIS software because of its widespread availability and complete set of processing tools. QGIS unfortunately lacks the ability to process certain proprietary data types from ArcGIS, the current industry standard, but is otherwise a capable and free option. Python, along with the GeoPandas package, is used to handle more complex modeling such as the carrying capacity and cost functions which would be difficult in GIS software alone. For these cases, a Python package is created to interface between the main model and the individual GIS data sets, as demonstrated in section 5.
@@ 4.1 Preprocessing data
Several preprocessing steps are necessary to standardize downloaded GIS data files and prepare them for further mapping and analysis. These include occasional file conversion to a QGIS-compatible format, vector interpolation for environmental conditions, clipping data sets to the study scope, and reprojection to either WGS 84 or WGS 84 / UTM Zone 19. More thorough instructions to complete these steps in QGIS are provided in A.
@@ 4.2 Raster analysis
Although most analysis for this case study is done within the Python model shown in section 5, there are several helpful functions in QGIS that can aid in understanding and visualizing the data. For this study, two are used: generating contours to display the range of oceanic conditions, and highlighting regions that meet certain conditions or constraints with Raster Calculator. Figure 4 shows all raster data sets used in the case study, including environmental conditions used to identify favorable conditions for a certain species, average wave period and height to estimate WEC power output, and bathymetry and distance to port which further constrain feasible sites and affect the cost. The environmental conditions are also used to evaluate the carrying capacity of a location, which indicates how large a population can be sustained in a given environment. Generating contours and using the Raster Calculator in QGIS are both discussed in A.
@@ 4.3 Vector analysis
In this study vector data is used to represent conflicts, which QGIS can aid in identifying. It is often helpful to convert between raster and vector formats depending on the analysis needed. Areas with high fishing vessel traffic, for example, are excluded in this analysis and so a vector of regions that have high fishing traffic is first created using the raster input data. Vector data sets are the core of the conflict analysis used to identify overlapping industries, and each of the included data sets are pictured in figure 5.
Like raster files, vector data sets are later loaded into Python, but preliminary maps of conflicts can also be generated in QGIS to identify and highlight potential issues early on. Figure 6 demonstrates this technique using the data sets for wind lease areas and regions of high fishing traffic, and highlights the regions where these industries overlap and therefore could conflict. Note that although the regions highlighted in figure 6 include areas that had above-average fishing traffic in 2021, only areas with very high fishing traffic (over one standard deviation above the average) are considered infeasible farm locations in the Python model. This conclusion was determined after attending meetings hosted by the Bureau of Ocean Energy Management with stakeholders in the fishing and offshore wind industries. The fishing community is very protective of their fishing grounds and information about where they are is limited. Therefore, this assumption reflects a reasonable understanding of potential conflicts with fisheries and other industries.